# reference.bib

### v1

@@ 1 Introduction
The intricacies and fluctuations of financial markets present substantial challenges for making high-quality, sequential investment decisions. In tasks such as single stock trading and portfolio management, each intelligent decision is informed by multiple market interactions and integrated information characterized by different levels of timeliness and modalities markowitz1952portfolio; dang2019squawk. These tasks aim to maximize profit while managing current market risks in an open-end environment.
@@
In practice, trading firms often rely on synthesized human teamwork, with an organizational structure that typically involves hierarchical communication among various functional roles such as data analysts, risk analysts, portfolio managers, etc., and careful integration of various resources maginn2007managing; radner1993organization. However, the cognitive limitations of human team members can restrict their ability to quickly process market signals and achieve optimal investment results miller1956magical.
@@
To improve investment returns and overcome human limitations, studies have employed methods such as deep reinforcement learning (DRL) to develop agent systems that simulate market environments and automate investment strategies wang2021commission; han2023select; qin2024earnhft. Meanwhile, advancements in large language models (LLMs) demonstrate potential in complex tasks such as reasoning, tool-using, planning, and decision-making song2023llm; paranjape2023art, promising to surpass existing agent architectures. Language agents are known for human-like communication and adaptable prompt-based structures tailored for diverse decision-making settings park2023generative; wu2023autogen; qiao2024autoact; hong2023metagpt. For optimal decision-making performance, two key elements need to be considered: 1) Organizing agents to promote effective teamwork and efficient communication, and 2) Enabling agents to continually learn and refine their actions. Studies indicate that imitating human organizational structures can effectively coordinate language agents for specific tasks guo2024embodied; li2024agent; chen2023agentverse. Additionally, recent developments in textual gradient-based prompt optimization pryzant2023automatic; tang2024unleashing and verbal reinforcement yao2023retroformer; shinn2024reflexion have proven to be practical for enhancing the reasoning and decision-making abilities of language agents iteratively.
@@
Language agent systems adapted for financial decision-making, such as FinGPT yang2023fingpt, FinMem yu2023finmem and FinAgent zhang2024finagent, have demonstrated impressive performance. However, these systems exhibit several limitations. Firstly, their reliance on the risk preferences of agents based on short-term market fluctuations fails to control overall risk exposures over long-term trading periods. This oversight can lead to potential losses due to the dismissal of fundamental factors driving investment profits. A more effective strategy is to quantify investment risks using measures of risk from quantitative finance delbaen2000coherent; fabozzi2010quantitative. Secondly, most existing systems are designed for single-asset trading tasks and show limited adaptability to other financial applications involving multiple assets, such as portfolio management. Thirdly, these systems depend on the information understanding and extraction ability of a single agent, placing heavy pressure on the language model to comprehend and process information within a limited context window. This likely degrades decision quality. A multi-agent system that allocates tasks based on data source and modality could enhance performance. Even though approaches like StockAgent liuai utilize a multi-agent system for individual stock trading tasks. However, their decision-making processes heavily depend on extensive discussions among numerous LLM agents, leading to high communication costs and extended processing times. Besides, the lack of a clear optimization goal in such systems can compromise the effectiveness of the outcomes.
@@
In response to these unaddressed issues, we propose FinCon, an LLM-based multi-agent framework for critical financial decision-making tasks such as single stock trading and portfolio management. Our main contributions are as follows: 1) Inspired by real-world investment roles, we designed an innovative Synthesized Manager-Analyst hierarchical communication structure and a risk control component. This communication structure distributes financial data from various sources and modalities to corresponding functional analyst agents, allowing them to focus on distilling essential investment insights and indicators from a single source of information. The manager then makes trading decisions by consolidating these insights and interacting with the risk-control component. The manager and analyst in a synchronous hierarchy to effectively enhance overall comprehension and reasoning capabilities. In this communication network, messages are conveyed only to the workers who need to be informed, eliminating redundant peer-to-peer communication and saving communication costs. 2) It is a generalized framework that can conduct not only stock trading but also portfolio management. The latter has not been claimed to be tackled by other financial language agent frameworks. 3) We innovatively crafted the agents’ persona using a dual-level risk control component, enabling comprehensive risk updates both within and across episodes. Within episodes, risk supervision is conducted using a quantile-based risk measure from quantitative finance, the Conditional Value at Risk (CVaR) kuester2006value. For across-episode risk control, we designed a unique verbal reinforcement method to update the decision maker’s investment beliefs with conceptualized investment insights to guide future decisions. These investment beliefs are distilled from underlying reasoning trajectory-level profit and loss (PnL) changes and summarized as conceptual perspectives based on the functionality of all analyst agents. Investment beliefs can then be selectively back-propagated from the manager to the corresponding functional analyst agent. Ablation study results illustrate the significant effectiveness of our risk control design in managing market risk and improving trading performance.
@@ 2 Related Work
LLM Agents for Financial Decision Making. There are considerable efforts directed towards developing general-purpose LLM agent for sequential decision-making shinn2023reflexion; zhao2024expel, and such type of tasks often involve episodic interactions with environment and verbal reflections for action refinement, such as coding competition li2022competition; chen2022codet, software development qian2023communicative; hong2023metagpt, game-playing xu2023language; ma2023large. Furthermore, researchers have started to exploit how LLM agents can perform better in harder decision-making tasks from finance de2023llm; zhao2024revolutionizing; cao2024risklabs, in which there are more volatile environments, leading to that the numerous unpredictable elements can obscure an agent’s ability to reflect accurately on the reasons for poor decision outcomes. FinMem yu2023finmem enhances single stock trading performance by embedding memory modules with LLM agent for reflection-refinement, and FinAgent zhang2024finagent improved trading profits via using external quantitative tool to fight against volatile environment.
Multi-Agent System and Communication Structures. In traditional multi-agent systems canese2021multi; zhang2021multi, the way for agents’ communication is pre-determined, like sharing data or state observations foerster2016learning; lin2021learning; kim2020communication; zhu2022survey; yao2024reinforcement. The emergence of large language model brings flexibility for human-understandable communications wang2023adapting; park2023generative; hong2023metagpt; mandi2023roco, so some work tries to elevate decision-making ability of LLM-based multi-agent system by letting agents engage in discussions zhang2023building; wu2023autogen or debates du2023improving; chan2023chateval. The similar peer-communication strategy was as well utilized by the multi-agent system for financial tasks xing2024designing; de2024optimized; wan2024enhancing. However, such approach are not optimal for unified-goal financial tasks that prioritize profits liuai2024, because they suffer from potentially ambiguous optimization objectives and are unable to control the unnecessary communication costs bolton1994firm.
Prompt Optimization and Verbal Reinforcement. To enhance the reasoning or decision-making of LLM agents, many prompt optimization techniques have been proposed, like ReAct yao2022react, Chain of Thought (CoT) wei2022chain, Tree of Thoughts (ToT) yao2024tree, ART paranjape2023art, intended for that LLM agents can automatically generate intermediate reasoning steps as an iterative program. In addition, to make LLM agent make decisions like humans and generate more understandable reasoning texts, some researchers recommend incorporating cognitive structures hao2023reasoning; zhu2023ghost; sumers2023cognitive; wang2024survey. Inspired by these previous work and DRL algorithms kaelbling1996reinforcement, verbal reinforcement yao2023retroformer; shinn2024reflexion; zhang2023controlling; guo2024embodied was developed for LLM agents such that they can update actions based on iterative self-reflection while integrating additional LLM as a prompt optimizer pryzant2023automatic; tang2024unleashing.
@@ 3 Preliminaries
Here, we outline the mathematical notations for the two major financial decision-making tasks that will be explicitly discussed in our work. We also formally present the generalized modeling formulation using a Partially Observable Markov Decision Process (POMDP) spaan2012partially for financial decision-making tasks.
@@ 3.1 Financial Decision-making Tasks Formulation
Portfolio Trading Tasks. In addition to processing multi-modal market information, the analyst agents also establish stocks’ pool for portfolio management, regarding statistical correlation between stocks’ returns. Then, the manager agent makes trading decisions on each stock in the pool. Lastly, the manager agent determines the portfolio weights for all stocks through a computational tool, which leverages the following mean-variance optimization fabozzi2008portfolio:
$\max_{\textbf{w}}\langle\textbf{w},\mu\rangle-\langle\textbf{w},\Sigma\textbf{w}\rangle\quad\text{s.t.}\leavevmode\nobreak\ w_{n}=\begin{cases}\in[0,1],&\text{``long''}\\ \in[-1,0],&\text{``short''}\\ =0,&\text{``neutral''}\end{cases},\leavevmode\nobreak\ \leavevmode\nobreak\ \forall n\in\{1,\cdots,N\}$
where $\textbf{w}=(w_{1}\cdots,w_{N})\in\mathbb{R}^{N}$ is portfolio weights vector, $\mu$ and $\Sigma$ are the shrinkage estimators of $N$-dimensional sample expected return and $N\times N$ sample covariance matrix of chosen stocks’ daily return sequences respectively fabozzi2010quantitative. We note that portfolio weights are rebalanced on daily basis. In our implementation, we first calculate the portfolio weights by solving the above optimization problem. Then, the target position is calculated by linearly scaling the portfolio weights from the prior step.
@@ 3.2 Modeling Quantitative Trading as POMDP
Then, our multi-agent system is supposed to learn the policies of all agents: the policies of analyst agents $\pi_{\theta^{i}}^{i}:\mathcal{X}\to\mathcal{A}^{i},i\in\{1,\cdots,I\}$ (the ways to process information, i.e. $\hat{O}_{t}^{i}\sim\pi_{\theta^{i}}^{i}(\cdot|O_{t}^{i})$), and the policy of manager agent $\pi_{\theta^{a}}:\mathcal{A}\times\mathcal{Y}\to\mathbb{A}$ (the ways to make trading decisions, i.e. $A_{t}\sim\pi_{\theta^{a}}(\cdot|\hat{O}_{t},B_{t})$) such that the system maximizes cumulative trading reward while controlling risk rao2022foundations.
@@ 3.3 Textual Gradient-Descent
In an LLM-based prompt optimizer, a meta-prompt pryzant2023automatic; tang2024unleashing is used to refine the task prompt for better performance. For example, for a mathematical reasoning task, the task prompt might be "Let’s solve the problem," while the meta-prompt could be "Improve the prompt to help a model better perform mathematical reasoning."
@@ 4.1.1 Manager-Analyst Agent Group
Analyst Agents. In FinCon, analyst agents distill concise investment insights from large volumes of multi-source market information for one specific trading target. To ensure high reasoning quality by reducing task load and enhancing analysis focus, each agent processes a single information source in a uni-modal manner, delivering pre-specified outputs based on prompts. This setup mirrors an effective human team structure, where each analyst specializes in a specific function and synchronously filters market noise to extract essential insights. These agents support the manager agent in making informed decisions by consolidating denoised investment information from multiple perspectives. We design seven distinct types of analyst agents using LLMs, each producing diverse investment insights, as shown in the upper part of Figure 2. Differentiated by input modalities, three textual data processing agents distill insights and sentiments from daily news and financial reports. An audio agent extracts investment tendencies from earnings call recordings using the Whisper API. Additionally, a data analysis agent and a stock selection agent compute key financial metrics, such as momentum and CVaR, based on tabular time series data. The stock selection agent also manages portfolio selection following the classic risk diversification method in quantitative finance markowitz1952portfolio, as detailed in Appendix. A.5.
@@ 4.1.2 Risk-Control Component
These investment belief updates are initially received by the manager agent and selectively propagated to relevant agent nodes, thereby minimizing over-communication. Additionally, unlike the text-based gradient descent proposed by Tang et al.tang2024unleashing, which uses prompt editing distance as the learning rate, we obtain investment belief prompt updates by measuring the overlapping percentage of trading actions between two adjacent training trajectories achieved at each investment belief update. This approach has proven effective in enhancing the performance of a synthesized agent system, where each worker has a clearly defined and specialized responsibility.
@@
     Compare two sets of conceptualized insights and give the reasoning for higher performance (providing textual optimization direction, i.e. $meta\leavevmode\nobreak\ prompt$);
@@
     Update the prompts by textual gradient-descent: $\bm{\theta}\longleftarrow M_{r}(\bm{\theta},\tau,meta\leavevmode\nobreak\ prompt).$
@@ 4.2 Modular Design of FinCon Agents
Here, we explain the modular design of FinCon agents. Inspired by the recent works of Park et al. sumers2023cognitive and Sumers et al. zhang2023exploring on developing the cognitive structure of language agents for human-like behavior, agents in FinCon integrate four modules to support their necessary functionalities, along with a shared general configuration, as detailed in Figure 3:
@@
Memory Module. The memory module contains three key components: working memory, procedural memory, and episodic memory. Similar to how humans process events in their working memory wagner1999working, FinCon agents utilize a working memory component to perform a series of operations, including observation, distillation, and refinement on available memory events, tailored to the agents’ specific roles. Procedural memory and episodic memory are essential for recording historical actions, outcomes, and reflections during sequential decision-making. Procedural memory is created after each decision step within an episode, storing data as memory events. For a trading inquiry, top events are retrieved from procedural memory, which are ranked based on recency, relevancy, and importance, which is the simplified version based on Yu et al. yu2023finmem and detailed in Appendix A.6. Different functional analyst agents possesses distinct procedural memory decay rates, reflecting the timeliness of various financial data sources, which is crucial for aligning multi-type data impacting a specific time point and facilitating informed decision-making. The manager agent augments the procedural memory of analyst agents by providing feedback through access counter. Procedural memory is maintained by both analyst and manager agents, each keeping different records, as shown in Figure 3. The episodic memory, exclusive to the manager agent, retains actions, PnL series of the previous episodes and conceptualized updated investment beliefs from the risk control component.

### v2

# FinCon: A Synthesized LLM Multi-Agent System with Conceptual Verbal Reinforcement for Enhanced Financial Decision Making

@@ 1 Introduction
The intricacies and fluctuations of financial markets present substantial challenges for making high-quality, sequential investment decisions. In tasks such as single stock trading and portfolio management, each intelligent decision is informed by multiple market interactions and integrated information characterized by different levels of timeliness and modalities [1, 2]. These tasks aim to maximize profit while managing current market risks in an open-end environment.
@@
In practice, trading firms often rely on synthesized human teamwork, with an organizational structure that typically involves hierarchical communication among various functional roles such as data analysts, risk analysts, portfolio managers, etc., and careful integration of various resources [3, 4]. However, the cognitive limitations of human team members can restrict their ability to quickly process market signals and achieve optimal investment results [5].
@@
To improve investment returns and overcome human limitations, studies have employed methods such as deep reinforcement learning (DRL) to develop agent systems that simulate market environments and automate investment strategies [6, 7, 8]. Meanwhile, advancements in large language models (LLMs) demonstrate potential in complex tasks such as reasoning, tool-using, planning, and decision-making [9, 10], promising to surpass existing agent architectures. Language agents are known for human-like communication and adaptable prompt-based structures tailored for diverse decision-making settings [11, 12, 13, 14]. For optimal decision-making performance, two key elements need to be considered: 1) Organizing agents to promote effective teamwork and efficient communication, and 2) Enabling agents to continually learn and refine their actions. Studies indicate that imitating human organizational structures can effectively coordinate language agents for specific tasks [15, 16, 17]. Additionally, recent developments in textual gradient-based prompt optimization [18, 19] and verbal reinforcement [20, 21] have proven to be practical for enhancing the reasoning and decision-making abilities of language agents iteratively.
@@
Language agent systems adapted for financial decision-making, such as FinGPT [22], FinMem [23] and FinAgent [24], have demonstrated impressive performance. However, these systems exhibit several limitations. Firstly, their reliance on the risk preferences of agents based on short-term market fluctuations fails to control overall risk exposures over long-term trading periods. This oversight can lead to potential losses due to the dismissal of fundamental factors driving investment profits. A more effective strategy is to quantify investment risks using measures of risk from quantitative finance [25, 26]. Secondly, most existing systems are designed for single-asset trading tasks and show limited adaptability to other financial applications involving multiple assets, such as portfolio management. Thirdly, these systems depend on the information understanding and extraction ability of a single agent, placing heavy pressure on the language model to comprehend and process information within a limited context window. This likely degrades decision quality. A multi-agent system that allocates tasks based on data source and modality could enhance performance. Even though approaches like StockAgent [27] utilize a multi-agent system for individual stock trading tasks. However, their decision-making processes heavily depend on extensive discussions among numerous LLM agents, leading to high communication costs and extended processing times. Besides, the lack of a clear optimization goal in such systems can compromise the effectiveness of the outcomes.
@@
In response to these unaddressed issues, we propose FinCon, an LLM-based multi-agent framework for critical financial decision-making tasks such as single stock trading and portfolio management. Our main contributions are as follows: 1) Inspired by real-world investment roles, we designed an innovative Synthesized Manager-Analyst hierarchical communication structure and a risk control component. This communication structure distributes financial data from various sources and modalities to corresponding functional analyst agents, allowing them to focus on distilling essential investment insights and indicators from a single source of information. The manager then makes trading decisions by consolidating these insights and interacting with the risk-control component. The manager and analyst in a synchronous hierarchy to effectively enhance overall comprehension and reasoning capabilities. In this communication network, messages are conveyed only to the workers who need to be informed, eliminating redundant peer-to-peer communication and saving communication costs. 2) It is a generalized framework that can conduct not only stock trading but also portfolio management. The latter has not been claimed to be tackled by other financial language agent frameworks. 3) We innovatively crafted the agents’ persona using a dual-level risk control component, enabling comprehensive risk updates both within and across episodes. Within episodes, risk supervision is conducted using a quantile-based risk measure from quantitative finance, the Conditional Value at Risk (CVaR) [28]. For across-episode risk control, we designed a unique verbal reinforcement method to update the decision maker’s investment beliefs with conceptualized investment insights to guide future decisions. These investment beliefs are distilled from underlying reasoning trajectory-level profit and loss (PnL) changes and summarized as conceptual perspectives based on the functionality of all analyst agents. Investment beliefs can then be selectively back-propagated from the manager to the corresponding functional analyst agent. Ablation study results illustrate the significant effectiveness of our risk control design in managing market risk and improving trading performance.
@@ 2 Related Work
LLM Agents for Financial Decision Making. There are considerable efforts directed towards developing general-purpose LLM agent for sequential decision-making [29, 30], and such type of tasks often involve episodic interactions with environment and verbal reflections for action refinement, such as coding competition [31, 32], software development [33, 14], game-playing [34, 35]. Furthermore, researchers have started to exploit how LLM agents can perform better in harder decision-making tasks from finance [36, 37, 38], in which there are more volatile environments, leading to that the numerous unpredictable elements can obscure an agent’s ability to reflect accurately on the reasons for poor decision outcomes. FinMem [23] enhances single stock trading performance by embedding memory modules with LLM agent for reflection-refinement, and FinAgent [24] improved trading profits via using external quantitative tool to fight against volatile environment.
Multi-Agent System and Communication Structures. In traditional multi-agent systems [39, 40], the way for agents’ communication is pre-determined, like sharing data or state observations [41, 42, 43, 44, 45]. The emergence of large language model brings flexibility for human-understandable communications [46, 11, 14, 47], so some work tries to elevate decision-making ability of LLM-based multi-agent system by letting agents engage in discussions [48, 12] or debates [49, 50]. The similar peer-communication strategy was as well utilized by the multi-agent system for financial tasks [51, 52, 53]. However, such approach are not optimal for unified-goal financial tasks that prioritize profits [54], because they suffer from potentially ambiguous optimization objectives and are unable to control the unnecessary communication costs [55].
Prompt Optimization and Verbal Reinforcement. To enhance the reasoning or decision-making of LLM agents, many prompt optimization techniques have been proposed, like ReAct [56], Chain of Thought (CoT) [57], Tree of Thoughts (ToT) [58], ART [10], intended for that LLM agents can automatically generate intermediate reasoning steps as an iterative program. In addition, to make LLM agent make decisions like humans and generate more understandable reasoning texts, some researchers recommend incorporating cognitive structures [59, 60, 61, 62]. Inspired by these previous work and DRL algorithms [63], verbal reinforcement [20, 21, 64, 15] was developed for LLM agents such that they can update actions based on iterative self-reflection while integrating additional LLM as a prompt optimizer [18, 19].
@@ 3 Preliminaries
Here, we outline the mathematical notations for the two major financial decision-making tasks that will be explicitly discussed in our work. We also formally present the generalized modeling formulation using a Partially Observable Markov Decision Process (POMDP) [65] for financial decision-making tasks.
@@ 3.1 Financial Decision-making Tasks Formulation
Portfolio Trading Tasks. In addition to processing multi-modal market information, the analyst agents also establish stocks’ pool for portfolio management, regarding statistical correlation between stocks’ returns. Then, the manager agent makes trading decisions on each stock in the pool. Lastly, the manager agent determines the portfolio weights for all stocks through a computational tool, which leverages the following mean-variance optimization [66]:
$\max_{\textbf{w}}\langle\textbf{w},\mu\rangle-\langle\textbf{w},\Sigma\textbf{w}\rangle\quad\text{s.t.}~{}w_{n}=\begin{cases}\in[0,1],&\text{``long''}\\ \in[-1,0],&\text{``short''}\\ =0,&\text{``neutral''}\end{cases},~{}~{}\forall n\in\{1,\cdots,N\}$
where $\textbf{w}=(w_{1}\cdots,w_{N})\in\mathbb{R}^{N}$ is portfolio weights vector, $\mu$ and $\Sigma$ are the shrinkage estimators of $N$-dimensional sample expected return and $N\times N$ sample covariance matrix of chosen stocks’ daily return sequences respectively [26]. We note that portfolio weights are rebalanced on daily basis. In our implementation, we first calculate the portfolio weights by solving the above optimization problem. Then, the target position is calculated by linearly scaling the portfolio weights from the prior step.
@@ 3.2 Modeling Quantitative Trading as POMDP
Then, our multi-agent system is supposed to learn the policies of all agents: the policies of analyst agents $\pi_{\theta^{i}}^{i}:\mathcal{X}\to\mathcal{A}^{i},i\in\{1,\cdots,I\}$ (the ways to process information, i.e. $\hat{O}_{t}^{i}\sim\pi_{\theta^{i}}^{i}(\cdot|O_{t}^{i})$), and the policy of manager agent $\pi_{\theta^{a}}:\mathcal{A}\times\mathcal{Y}\to\mathbb{A}$ (the ways to make trading decisions, i.e. $A_{t}\sim\pi_{\theta^{a}}(\cdot|\hat{O}_{t},B_{t})$) such that the system maximizes cumulative trading reward while controlling risk [70].
@@ 3.3 Textual Gradient-Descent
In an LLM-based prompt optimizer, a meta-prompt [18, 19] is used to refine the task prompt for better performance. For example, for a mathematical reasoning task, the task prompt might be "Let’s solve the problem," while the meta-prompt could be "Improve the prompt to help a model better perform mathematical reasoning."
@@ 4.1.1 Manager-Analyst Agent Group
Analyst Agents. In FinCon, analyst agents distill concise investment insights from large volumes of multi-source market information for one specific trading target. To ensure high reasoning quality by reducing task load and enhancing analysis focus, each agent processes a single information source in a uni-modal manner, delivering pre-specified outputs based on prompts. This setup mirrors an effective human team structure, where each analyst specializes in a specific function and synchronously filters market noise to extract essential insights. These agents support the manager agent in making informed decisions by consolidating denoised investment information from multiple perspectives. We design seven distinct types of analyst agents using LLMs, each producing diverse investment insights, as shown in the upper part of Figure 2. Differentiated by input modalities, three textual data processing agents distill insights and sentiments from daily news and financial reports. An audio agent extracts investment tendencies from earnings call recordings using the Whisper API. Additionally, a data analysis agent and a stock selection agent compute key financial metrics, such as momentum and CVaR, based on tabular time series data. The stock selection agent also manages portfolio selection following the classic risk diversification method in quantitative finance [1], as detailed in Appendix. 7.5.
@@ 4.1.2 Risk-Control Component
These investment belief updates are initially received by the manager agent and selectively propagated to relevant agent nodes, thereby minimizing over-communication. Additionally, unlike the text-based gradient descent proposed by Tang et al.[19], which uses prompt editing distance as the learning rate, we obtain investment belief prompt updates by measuring the overlapping percentage of trading actions between two adjacent training trajectories achieved at each investment belief update. This approach has proven effective in enhancing the performance of a synthesized agent system, where each worker has a clearly defined and specialized responsibility.
@@
     Compare two sets of conceptualized insights and give the reasoning for higher performance (providing textual optimization direction, i.e. $meta~{}prompt$);
@@
     Update the prompts by textual gradient-descent: $\bm{\theta}\longleftarrow M_{r}(\bm{\theta},\tau,meta~{}prompt).$
@@ 4.2 Modular Design of FinCon Agents
Here, we explain the modular design of FinCon agents. Inspired by the recent works of Park et al. [61] and Sumers et al. [71] on developing the cognitive structure of language agents for human-like behavior, agents in FinCon integrate four modules to support their necessary functionalities, along with a shared general configuration, as detailed in Figure 3:
@@ 1. Working:        - Observation    - Retrieval    - Distillation
Memory Module. The memory module contains three key components: working memory, procedural memory, and episodic memory. Similar to how humans process events in their working memory [72], FinCon agents utilize a working memory component to perform a series of operations, including observation, distillation, and refinement on available memory events, tailored to the agents’ specific roles. Procedural memory and episodic memory are essential for recording historical actions, outcomes, and reflections during sequential decision-making. Procedural memory is created after each decision step within an episode, storing data as memory events. For a trading inquiry, top events are retrieved from procedural memory, which are ranked based on recency, relevancy, and importance, which is the simplified version based on Yu et al. [23] and detailed in Appendix 7.6. Different functional analyst agents possesses distinct procedural memory decay rates, reflecting the timeliness of various financial data sources, which is crucial for aligning multi-type data impacting a specific time point and facilitating informed decision-making. The manager agent augments the procedural memory of analyst agents by providing feedback through access counter. Procedural memory is maintained by both analyst and manager agents, each keeping different records, as shown in Figure 3. The episodic memory, exclusive to the manager agent, retains actions, PnL series of the previous episodes and conceptualized updated investment beliefs from the risk control component.